# Mitigating Text Toxicity with Counterfactual Generation

### v1

@@ 5 On the use and misuse of toxicity detection and mitigation tools
Online toxicity is a systemic problem with complex and multiple roots Salminen et al. (2018). Automatic toxicity detection and mitigation do not in themselves solve the factors causing online toxic content generation, but they offer technological means of adaptation to its rapid online development. The use of such tools raises critical ethical and technical considerations. In this section we identify some of the risks associated with the use of toxicity mitigation tools, and propose good practice rules to limit these risks.
5.1 Diversity of values and control of information
In this work, we have considered a usual hate speech definition as "aggressive or offensive language that can be focused on a specific group of people who share common property such as religion, race, gender, sexual orientation, sex or political affiliation". This definition is just one of many used by institutions and platforms to characterize hate speech Fortuna and Nunes (2018). In particular, hate speech characterization can focus either on violence and hate incentives or on the objective of directly attacking. The choice of a specific definition can have a direct impact on the way online toxicity is automatically processed with automatic toxicity processing tools.
@@ 5.1 Diversity of values and control of information
In addition, the perceived toxicity of language can vary based on identity and beliefs Al Kuwatly et al. (2020). For instance, conservative annotators can show a higher propensity to label African American English dialect as toxic while being less likely to annotate anti-Black comments as harmful Sap et al. (2022). Such annotator bias can be reflected in the datasets used to detect toxicity. Recommendation No. 1: When building databases to train ML models to detect toxicity, data annotators must be selected to represent this diversity of values. Besides, toxicity mitigation with NLP must be done by carefully selecting the toxicity dataset used to train a toxicity classifier, in order to make sure that the values implicitly encoded in the classifier match those expected.
Toxicity mitigation is less restrictive than full detection and removal of toxic messages, since the non-toxic part of the initial text is supposed to be retained. However, having a comment moderated by a toxicity mitigation algorithm can be perceived as a censorship and unjustified information control mechanism. Recommendation No. 2: The toxicity definitions chosen and the datasets used to train the toxicity detector have to be made transparently accessible. Making this information transparent and easily accessible is more likely to ensure a relationship of trust between the content generator and the platform on which the text is posted. The vision chosen to define toxicity could then be subject to deliberative questioning by the users.
@@ 5.2 Against malicious use
A toxicity mitigation tool can be misused in various ways. Since toxicity mitigation implies to learn how to detect toxic content and replace it, the same process can be carried out in the opposite way to poison text. This way, counterfactually generating toxicity would consist in starting from a non-toxic text and slightly modifying it to make it toxic using the same toxicity classifier used to perform toxicity mitigation. Such a way of using the tool would automatically turn non-toxic texts into more toxic ones while still being plausible. Recommendation No. 3: One way of preventing counterfactual toxic content generation is to fine-tune the neural models performing text modification on detoxified text corpora. A toxicity vocabulary can also be used to prevent the language model from generating text within it. However, these solutions cannot provide a complete guarantee against someone using the tool as a toxicity amplifier to poison text.
@@
Another misuse of a toxicity mitigation tool is to use it as an adversarial attack generator to make a toxic text seem non-toxic. Adversarial attacks are small perturbations of data instances fooling a classifier with imperceptible changes Liang et al. (2022), which bring them formally close to counterfactuals. Since toxicity mitigation methods are based on the use of a toxicity classifier to steer the detoxification process, these methods are subject to adversarial attacks. In this manner, a dishonest user could hijack a detoxification tool to find the smallest modifications to the initial texts, leading a toxicity classifier to falsely assess that a text is correctly detoxified. Recommendation No. 4: A recent work Colombo et al. (2023) proposes a method to robustly detect textual adversarial attacks based on the computation of similarities between a given input embedding and the training distribution. Another way to prevent adversarial attacks is to make the toxicity classifier more robust through counterfactual data augmentation Wu et al. (2021).
Toxicity mitigation methods can also be reverse engineered to discover the rules used by an online platform to detect and modify harmful contents. This can lead to a change in the terms used, expressed in a seemingly neutral way in order to continue to publicly expressing hateful content online Ji and Knight (2018). Toxicity classifiers have to be frequently updated and fine-tuned on updated datasets integrating changes in the vocabulary used to express toxic ideas.
5.3 The inaccuracy of automatic toxicity processing
State-of-the-art toxicity mitigation algorithms do not remove toxicity with perfect precision. Therefore, deploying a toxicity mitigation algorithm fully automatically is particularly risky since it could let harmful content spread. We propose to use this kind of tool as the first layer of hate content processing before integrating humans into the loop.
@@ 5.3 The inaccuracy of automatic toxicity processing
Behind any platform, a lot of content must be reviewed and online moderation is partly performed by human labor Gillespie (2018); Jhaver et al. (2023). By being exposed to disturbing toxic content, human moderators can develop psychological and emotional distress Spence et al. (2023). Toxicity mitigation tools have the potential to induce a socio-technical change, suggesting textual intervention and reducing exposure for content moderators. Recommendation No. 5: We suggest using hate content detectors and mitigation methods based on the level of toxicity of the text. Text with the highest level of toxicity could simply be deleted, as it would be unlikely to be modified without completely altering its original meaning. Intermediate toxicity levels could be handled by a toxicity mitigation algorithm in order to preserve the general meaning of the text while proposing a softened version. This way, content moderators’ exposure to the most hateful content would be significantly limited, and the more ambiguous content would be preprocessed by the mitigation algorithm to propose a more acceptable first version, still requiring moderation.
@@ 7 Conclusion
8 Ethic statement
Each participant to the human evaluation signed an informed consent form outlining the project’s purpose and details, and the intended use of the data they would generate. The data were anonymized and processed only by the authors. The data produced are stored in a file in accordance with the General Data Protection Regulation (GDPR) regulations in force. Participation in the study was fully voluntary. It was possible to stop performing the labeling tasks at any time. The Consent form used is anonymized and presented in the Appendix, Figure 6. The authors of this paper do not represent any organization or institution engaged in data labeling activities. This study was conducted for research purposes only.
For anonymity reasons, we do not provide a link to the GitHub of the paper. The code will be available upon acceptance to facilitate reproduction and further research.

### v2

@@ 5 Discussion
Online toxicity is a systemic problem with complex and multiple roots Salminen et al. (2018). Automatic toxicity processing does not in itself solve the factors causing online toxic content generation, but they offer technological means of adaptation to its rapid development. While CF toxicity mitigation may yield competitive results, it also raises ethical concerns that we discuss in Section 7.
6 Conclusion
This paper formalized how LFI and CF generation methods and CFI can be leveraged to accurately target textual toxic content and perform toxicity mitigation. CF-Detox leads to competitive results, with state-of-the-art performance in terms of content preservation while accurately detoxifying text and generating plausible text. CF-Detox is versatile since it can be used with various types of LFI methods (such as attention, gradient and perturbation) to target toxic POS and any kind of classifier allowing LFI computation.
This paper is the first to show the extent to which fields such as automatic toxicity processing and explainable AI, which have developed in parallel, actually share epistemic similarities and can be mutually beneficial, paving the way towards practical applications of XAI methods.
7 Ethical Considerations
The use of automatic processing tools raises crucial ethical considerations. We identify some of the risks associated with the use of toxicity mitigation tools, and draw recommendations to limit them.
7.1 Managing Diversity of Values
@@
In this work, we have considered a usual hate speech definition as "aggressive or offensive language that can be focused on a specific group of people who share common property such as religion, race, gender, sexual orientation, sex or political affiliation". This definition is just one of many used by institutions and platforms to characterize hate speech Fortuna and Nunes (2018). In particular, hate speech characterization can focus either on violence and hate incentives or on the objective of directly attacking. The choice of a specific definition can have a tangible impact on the way online toxicity is automatically processed and perceived by web users. In addition, the perceived toxicity of language can vary based on identity and beliefs Al Kuwatly et al. (2020). For instance, conservative annotators can show a higher propensity to label African American English dialect as toxic while being less likely to annotate anti-Black comments as harmful Sap et al. (2022). Such annotator biases can be reflected in the datasets used to detect toxicity. Recommendation: Toxicity training sets have to be built by annotators carefully selected to represent diversity of values. Besides, toxicity training sets have to be chosen in order to to make sure that the values implicitly encoded in the classifier match those expected.
@@ 7.2 Against Malicious Use
A toxicity mitigation tool can be misused in various ways. One example is to use it as an adversarial attack generator to make a toxic text seem non-toxic. Adversarial attacks are small perturbations of data instances fooling a classifier with imperceptible changes Liang et al. (2022), which bring them formally close to CFs. Since toxicity mitigation methods are based on the use of a toxicity classifier to steer the detoxification process, these methods are subject to adversarial attacks. In this manner, a dishonest user could hijack a detoxification tool to find small modifications to the initial texts, leading a toxicity classifier to falsely assess that a text is correctly detoxified. Recommendation: A recent work Colombo et al. (2023) proposes a method to detect textual adversarial attacks based on the computation of similarities between an input embedding and the training distribution. Toxicity mitigation methods can also be reverse engineered to discover the rules used by an online platform to process harmful contents. This can lead to a change in the terms used, expressed in a seemingly neutral way in order to continue to express hateful content online Ji and Knight (2018). Recommendation: Toxicity classifiers have to be frequently retrained on updated datasets integrating changes in the vocabulary used to express toxicity to overcome toxicity concept drift.
7.3 Favoring Human-in-the-loop to Overcome Detoxification Inaccuracy
@@
State-of-the-art toxicity mitigation algorithms do not remove toxicity perfectly. Therefore, deploying a toxicity mitigation algorithm fully automatically is particularly risky since it could let harmful content spread. We propose to use these tools as the first layer of hate content processing before integrating humans into the loop. Behind any platform, online content must be reviewed and moderation is partly performed by human labor Gillespie (2018); Jhaver et al. (2023). By being exposed to disturbing toxic content, human moderators can develop psychological and emotional distress Spence et al. (2023). Toxicity mitigation tools have the potential to induce a socio-technical change, suggesting textual intervention and reducing exposure for content moderators. Recommendation: We suggest using hate content detectors and mitigation methods based on the level of toxicity of the text. Text with the highest level of toxicity could simply be deleted, as it would be unlikely to be modified without completely altering its original meaning. Intermediate toxicity levels could be handled by a toxicity mitigation algorithm in order to preserve the general meaning of the text while proposing a softened version. This way, content moderators’ exposure to the most hateful content would be significantly limited, and the more ambiguous content would be preprocessed by the mitigation algorithm to propose a more acceptable first version.
7.4 Personal Data
Each voluntary participant to the human evaluation signed an informed consent form outlining the project’s purpose and details. The data were anonymized and processed only by the authors and stored in accordance with the General Data Protection Regulation (GDPR). It was possible to stop at any time. The Consent form used is anonymized and presented in the supplementary material. This study was conducted for research purposes only.